# LLM3: Large Language Model-based Task and Motion Planning with Motion Failure Reasoning

### v2

@@ Abstract
Conventional Task and Motion Planning (TAMP) approaches rely on manually crafted interfaces connecting symbolic task planning with continuous motion generation. These domain-specific and labor-intensive modules are limited in addressing emerging tasks in real-world settings. Here, we present LLM3, a novel Large Language Model (LLM)-based TAMP framework featuring a domain-independent interface. Specifically, we leverage the powerful reasoning and planning capabilities of pre-trained LLMs to propose symbolic action sequences and select continuous action parameters for motion planning. Crucially, LLM3 incorporates motion planning feedback through prompting, allowing the LLM to iteratively refine its proposals by reasoning about motion failure. Consequently, LLM3 interfaces between task planning and motion planning, alleviating the intricate design process of handling domain-specific messages between them. Through a series of simulations in a box-packing domain, we quantitatively demonstrate the effectiveness of LLM3 in solving TAMP problems and the efficiency in selecting action parameters. Ablation studies underscore the significant contribution of motion failure reasoning to the success of LLM3. Furthermore, we conduct qualitative experiments on a physical manipulator, demonstrating the practical applicability of our approach in real-world settings.
@@ I Introduction
Sequential manipulation planning is an essential capability for robots to autonomously perform diverse tasks in complex environments. Executable motions must be effectively generated for robots to achieve long-term task objectives, requiring efficient planning algorithms for responsive operation and reasoning capabilities to anticipate environmental changes. Task and Motion Planning (TAMP) formulates a promising methodology that hierarchically decomposes planning into two stages: the high-level symbolic task planning stage reasons over long-horizon abstract action sequences, and the low-level continuous motion planning stage computes feasible trajectories subject to geometric constraints. In recent years, TAMP has enabled significant advances in diverse applications [1, 2, 3, 4, 5, 6]. However, a persistent challenge remains to properly interface between the task planner and the motion planner to efficiently solve TAMP, i.e., generating action sequences that satisfy both symbolic task goals and continuous motion constraints.
@@
Recent Large Language Models (LLMs[10] pre-trained on web-scale text data have demonstrated emergent capabilities in reasoning [11] and planning [12] when provided with in-context prompts [13, 14]. Pre-trained LLMs can perform task planning [12, 15, 16], generate continuous parameters [17], and reason on environment feedback [18, 16]. Our intuition is that pre-trained LLMs could provide a general and domain-independent approach to interfacing between symbolic and continuous domains for TAMP, eliminating the need to design domain-specific modules manually.
@@ I-A Related Work
Traditional TAMP approaches employ a high-level task planner to generate symbolic action sequences and a low-level motion planner to generate motion trajectories. The task planner requires pre-designed symbolic planning domains represented in formatted representations, such as Planning Domain Definition Language (PDDL). Significant efforts have been made to develop manually engineered modules that interface the task planner and motion planner, such as incorporating motion-level constraints into task planning [19, 4], making approximations at the motion level [20, 21], and designing specialized communication modules [9]. However, manually defining task planning domains and interface modules to fully capture real-world complexity is impractical. Furthermore, as the action space grows, searching for geometrically feasible symbolic action sequences becomes computationally challenging without effective heuristics [19]. Recent work has explored data-driven heuristics to improve TAMP efficiency [7, 22, 23, 24], but these domain-specific heuristics lack generalizability across domains. In this work, we employ a pre-trained LLM as both the task planner and the interface between task and motion. We expect that semantic knowledge in the LLM can provide domain-independent heuristics for TAMP.
@@
Recent Large Language Models (LLMs[10] encode vast world knowledge and exibit the emergent capability for planning [12, 25] through few-shot or zero-shot in-context learning [13, 14]. Pre-trained LLMs have been applied for task planning of robots or embodied agents [15, 16, 26, 27, 28, 29, 18, 30, 31]. Notably, Inner Monologue [16] takes in textualized environment feedback and generate actions to execute, while ReAct [18] further advanced this closed-loop approach by integrating reasoning and acting. Voyager [29] and DEPS [28] focus on developing open-ended embodied agents that iteratively replan based on execution failure in PC games. Furthermore, LLMs have been applied to solve TAMP problems  [32, 33]. Specifically, LLM-GROP [32] employs a pre-trained LLM to instantiate symbolic goals and continuous object placements for semantic object rearrangement, which are then fed into a classical TAMP planner. AutoTAMP [33] leverages an LLM to translate natural-language task specifications into a formal language processable by off-the-shelf TAMP algorithms. Our usage of LLMs in TAMP is inspired by many of the above works; however, the major difference is that we leverage the LLM as the core component of our TAMP framework.
I-B Overview
The remainder of this paper is organized as follows. We first provide preliminaries of TAMP and LLM-based planning in Section II. Then we elaborate on the proposed LLM3 framework in Section III. We summarize the results of simulated and real robot experiments and provide a discussion in Section IV. Finally, Section V concludes the paper and acknowledges the limitations of LLM3.
Figure 2: The system diagram of the proposed LLM3 framework. (a) We show an example of utilizing a pre-trained LLM for reasoning and generating action sequences. (b) The feasibility of the proposed action sequence is verified by rollout with a motion planner and transition function $\mathcal{T}$. The motion planning feedback is saved into a trace that is provided to the LLM in the next iteration.
@@ II-A Task and Motion Planning
A Task and Motion Planning (TAMP) problem [3] is a tuple $<\mathcal{O},\mathcal{S},\mathcal{A},\mathcal{T},s_{0},g>$, where:
@@
A common strategy to solve the TAMP problem is known as “search-then-sample” [9, 1, 3], which alternates generating symbolic action sequences through backtracking search and sampling the continuous action parameters until a feasible plan is found to reach the goal. However, a naive search-then-sample task and motion planner is usually inefficient, as verifying the action feasibility requires invoking the computationally expensive motion planning process. To mitigate this complexity, researchers have crafted symbolic domains [9, 34] to prune infeasible symbolic action sequences, or learned heuristic samplers to sample action parameters [8, 35] that lead to feasible plans. In this work, we leverage pre-trained LLMs as both a task planner and a heuristic sampler to generate symbolic action sequences and parameters that tend to produce feasible plans.
@@ II-B Planning as Sequence Prediction
Recently, planning problems have been formulated as sequence prediction to avoid the computationally expensive search process [36, 37, 38]. In particular, pre-trained LLMs are employed to generate discrete actions [12, 25] and continuous parameters [26] in an auto-regressive manner when provided in-context prompts [14]. Formally, with textualized initial state $s_{0}$, goal $g$ and additional context $c$, we have:
@@ III Method
We introduce LLM3, a TAMP framework that leverages a pre-trained LLM to reason on motion failure and generate iteratively refined symbolic actions and continuous action parameters. The system diagram of LLM3 is shown in Fig. 2. Below, we elaborate on the overall LLM3 framework, reasoning and planning with pre-trained LLM, and the designed motion planning feedback.
@@
// initialize output and feedback trace $plan\leftarrow[]$, $success\leftarrow\text{False}$, $feedback\leftarrow[]$, $trace\leftarrow[]$ // initialize current state and planning iteration count $s\leftarrow s_{0}$, $iter\leftarrow 0$ // main planning loop while not $success$ and $iter<N_{max}$ do
       // reason and plan with LLM $reason,llm\_plan\leftarrow LLM(s_{0},trace)$, $iter\leftarrow iter+1$ // rollout $llm\_plan$ from $s$ with motion planner foreach $\underline{a}\in llm\_plan$ do
             // call motion planner to verify feasibility of a $\tau,mp\_feedback\leftarrow MP(s;\underline{a})$ $feedback.append((\underline{a},mp\_feedback))$ // terminate rollout if action $\underline{a}$ is infeasible if $\tau$ is $None$ then
@@
            // update state $s$ and add $\underline{a}$ to $plan$ $s\leftarrow\mathcal{T}(s,\underline{a}(\tau))$, $plan.append(\underline{a}(\tau))$
      // check whether $s$ satisfies goal $success,task\_feedback\leftarrow g(s)$ // when the plan is feasible but doesn’t reach goal if not $success$ then
             // record $task\_feedback$ $feedback.append(task\_feedback)$ // add aggregated feedback to $trace$ $trace.append(feedback)$, $trace\leftarrow trace[-k:]$ // clear aggregated feedback $feedback\leftarrow[]$
2
return $success$, $plan$
@@ III-B Reasoning and Planning with pre-trained LLM
Following previous attempts in utilizing LLMs for reasoning and planning [11, 12, 16, 18], we prompt a pre-trained LLM to generate motion failure reasoning and action sequences in text format. Since we want to limit the domain-specific prior provided to the LLM, we use zero-shot prompting [11] without providing any planning examples. We adopt Chain-of-Thought (CoT) prompting [39, 18] to have the LLM generate the reasoning and action sequence in an auto-regressive manner. The prompt fed into LLM comprises the following contents: (i) a system message that provides the global context to the pre-trained LLM and activates its planning capability, (ii) a task description specifying the environment, goal, and available primitive actions, (iii) the textualized initial environment state, (iv) the trace of motion planning feedback, and (v) the output format. Note that most promp content will remain unchanged for different planning iterations of the same TAMP problem. Only the motion planning feedback trace in (iv) will be updated as new failed action sequences are added. Fig. 2(a) illustrated an example of reasoning and planning with LLM.
@@
[…]
@@ III-C Synthesizing Motion Planning Feedback
    The goal configuration has no feasible inverse kinematics (IK) solution.
@@ IV-A Simulation Setup
We developed a PyBullet-based simulation environment for our box-packing tasks, as illustrated in Fig. 5a. In Setting 1, three different sets of objects are given with increasing total sizes, while the size of the basket remains constant. This task requires the LLM3 to oversee potential collisions between objects or the robot throughout the action sequence. The LLM must reason why collisions occur and adjust previous actions to ensure feasible task and motion plans. Setting 2 involves placing the Set 3 objects into baskets of increasing sizes. Here, the robot cannot access the entire basket region but encounters a collision likelihood similar to the most crowded condition in Setting 1. In this setting, the task becomes more complex as it challenges the LLM to reason about the robot’s operational space and adjust the plan accordingly. Throughout the simulations, we utilize GPT-4 Turbo [41] as the LLM planner and BiRRT [40] as the motion planner, with 10 attempts for each setting. The code for replicating the experiment is available at https://github.com/AssassinWS/LLM-TAMP.
@@ IV-B Ablation Study
    LLM3 Scratch: The proposed LLM3 framework from scratch variant. It replans the entire action sequence if any action fails, incorporating motion planning feedback. See Fig. 3.
@@
In both settings, integrating motion planning feedback results in a decrease in #LM and #MP, along with an increase in %SR for both backtrack and scratch strategies. This indicates that the LLM can reason about failures from motion planning feedback, and importantly, propose adjusted task plans and action parameters that are more likely to produce feasible motions. Surprisingly, there is no clear evidence to suggest that utilizing backtracking is superior to replanning from scratch. We’ve observed distinct behaviors among different strategies employed by LLM3. Specifically, when utilizing backtracking, LLM sometimes consistently adjusts the action parameter of the specific action that directly led to failure, without adjusting actions several steps prior. In contrast, the LLM3 employing a planning from scratch strategy simply re-samples all actions and parameters, occasionally resulting in a slightly lower #LM.
@@ IV-C Action Parameter Selection
Random Sampling 109.6 663.1
LLM Sampling 10.8 70.2
LLM Sampling + Feedback 7.9 53.2
@@ IV-D Experiment Setup
To validate the effectiveness of our proposed method in a real-world setting, we conducted an experiment using a Franka Research 3 manipulator. The goal was to demonstrate the robot’s ability to perceive and manipulate objects in an environment with uncertainties in both perception and execution. The robot observed a single point cloud from a third-person-view RGB-D camera (Kinect Azure), capturing the workspace containing various objects such as blocks and a plate. To identify and locate individual objects, we employed Grounded Segment Anything (Grounded-SAM) [42] for object segmentation, initially segmenting objects in the 2D RGB image and then projecting the results onto the corresponding 3D point cloud. This approach yielded per-object point clouds, essential for planning and executing manipulation tasks.

### v3

@@ Abstract
Conventional Task and Motion Planning (TAMP) approaches rely on manually designed interfaces connecting symbolic task planning with continuous motion generation. These domain-specific and labor-intensive modules are limited in addressing emerging tasks in real-world settings. Here, we present LLM3, a novel Large Language Model (LLM)-based TAMP framework featuring a domain-independent interface. Specifically, we leverage the powerful reasoning and planning capabilities of pre-trained LLMs to propose symbolic action sequences and select continuous action parameters for motion planning. Crucially, LLM3 incorporates motion planning feedback through prompting, allowing the LLM to iteratively refine its proposals by reasoning about motion failure. Consequently, LLM3 interfaces between task planning and motion planning, alleviating the intricate design process of handling domain-specific messages between them. Through a series of simulations in a box-packing domain, we quantitatively demonstrate the effectiveness of LLM3 in solving TAMP problems and the efficiency in selecting action parameters. Ablation studies underscore the significant contribution of motion failure reasoning to the success of LLM3. Furthermore, we conduct qualitative experiments on a physical manipulator, demonstrating the practical applicability of our approach in real-world settings.
@@ I Introduction
Sequential manipulation planning is an essential capability for robots to autonomously perform diverse tasks in complex environments. Executable motions must be effectively generated for robots to achieve long-term task objectives, requiring efficient planning algorithms for responsive operation and reasoning capabilities to anticipate environmental changes. Task and Motion Planning (TAMP) formulates a methodology that hierarchically decomposes planning into two stages: the high-level symbolic task planning stage reasons over long-horizon abstract action sequences, and the low-level continuous motion planning stage computes feasible trajectories subject to geometric constraints. In recent years, TAMP has enabled significant advances [1, 2, 3, 4, 5, 6, 7]. However, a persistent challenge remains to properly interface between the task planner and the motion planner to efficiently solve TAMP, i.e., generating action sequences that satisfy both symbolic task goals and continuous motion constraints.
@@
Recent Large Language Models (LLMs) pre-trained on web-scale text data have demonstrated emergent capabilities in reasoning [11] and planning [12]. Pre-trained LLMs can perform task planning [13], generate continuous parameters [14], and reason on environment feedback [13]. Our intuition is that pre-trained LLMs could provide a general and domain-independent approach to interfacing between symbolic and continuous domains for TAMP, eliminating the need to design domain-specific modules manually.
@@ I-A Related Work
Conventional TAMP approaches employ a high-level task planner to generate symbolic action sequences and a low-level motion planner to generate motion trajectories. The task planner requires pre-designed symbolic planning domains represented in formatted representations, such as Planning Domain Definition Language (PDDL). Significant efforts have been made to develop manually engineered modules that interface the task planner and motion planner, such as incorporating motion-level constraints into task planning [15, 5], making approximations at the motion level [16, 17], and designing specialized communication modules [10]. However, manually defining task planning domains and interface modules to fully capture real-world complexity is impractical. Furthermore, as the action space grows, searching for geometrically feasible symbolic action sequences becomes computationally challenging without effective heuristics [15]. Recent work has explored data-driven heuristics to improve TAMP efficiency [8, 18], but these domain-specific heuristics lack generalizability across domains. In this work, we employ a pre-trained LLM as both the task planner and the interface between task and motion. We expect that semantic knowledge in the LLM can provide domain-independent heuristics for TAMP.
@@
Recent Large Language Models (LLMs[19, 20] encode vast world knowledge and exibit the emergent capability for planning [12]. Pre-trained LLMs have been applied for task planning of robots or embodied agents [13, 21, 22, 23, 24, 25, 26, 27]. Notably, Inner Monologue [13] takes in textualized environment feedback and generate actions to execute, while ReAct [28] further advanced this closed-loop approach by integrating reasoning and acting. Voyager [22] and DEPS [21] focus on developing open-ended embodied agents that iteratively replan based on execution failure. Furthermore, LLMs have been applied to solve TAMP problems  [29, 30]. Specifically, LLM-GROP [29] employs a pre-trained LLM to instantiate symbolic goals and continuous object placements for semantic object rearrangement, which are then fed into a classical TAMP planner. AutoTAMP [30] leverages an LLM to translate natural-language task specifications into a formal language processable by off-the-shelf TAMP algorithms. Our usage of LLMs in TAMP is inspired by many of the above works; however, the major difference is that we leverage the LLM as the core component of our TAMP framework.
Figure 2: System diagram of the proposed LLM3 framework. (a) We show an example of utilizing a pre-trained LLM for reasoning and generating action sequences. (b) The feasibility of the proposed action sequence is verified by rollout with a motion planner and transition function $\mathcal{T}$. The motion planning feedback is saved into a trace that is provided to the LLM in the next iteration.
@@ II-A Task and Motion Planning
A TAMP problem is a tuple $<\mathcal{O},\mathcal{S},\mathcal{A},\mathcal{T},s_{0},g>$, where:
@@
A common strategy to solve the TAMP problem is known as “search-then-sample” [10, 1, 3], which alternates generating symbolic action sequences through backtracking search and sampling the continuous action parameters until a feasible plan is found to reach the goal. However, a naive search-then-sample task and motion planner is usually inefficient, as verifying the action feasibility requires invoking the computationally expensive motion planning process [31, 32]. To mitigate this complexity, researchers have crafted symbolic domains [10, 33, 34] to prune infeasible symbolic action sequences, or learned heuristic samplers to sample action parameters [9, 35] that lead to feasible plans. In this work, we leverage pre-trained LLMs as both a task planner and a heuristic sampler to generate symbolic action sequences and action parameters.
@@ II-B Planning as Sequence Prediction
Recently, planning problems have been formulated as sequence prediction to avoid the computationally expensive search process [36]. In particular, pre-trained LLMs are employed to generate discrete actions [12] and continuous parameters in an auto-regressive manner when provided in-context prompts. Formally, with textualized initial state $s_{0}$, goal $g$ and additional context $c$, we have:
@@ III Method
We introduce LLM3, a TAMP framework that leverages a pre-trained LLM to reason on motion failure and generate iteratively refined symbolic actions and continuous action parameters. The system diagram of LLM3 is shown in Fig. 2. Below, we elaborate on the overall framework, reasoning and planning with the pre-trained LLM, and the designed motion planning feedback.
1
@@
2
// initialize output and feedback trace
3 $plan\leftarrow[]$, $success\leftarrow\text{False}$, $feedback\leftarrow[]$, $trace\leftarrow[]$
// initialize current state and planning iteration count
4 $s\leftarrow s_{0}$, $iter\leftarrow 0$
// main planning loop
5 while not $success$ and $iter<N_{max}$ do
       // reason and plan with LLM
6       $reason,llm\_plan\leftarrow LLM(s_{0},trace)$, $iter\leftarrow iter+1$
       // rollout $llm\_plan$ from $s$ with motion planner
7       foreach $\underline{a}\in llm\_plan$ do
             // call motion planner to verify feasibility of a
8             $\tau,mp\_feedback\leftarrow MP(s;\underline{a})$
9             $feedback.append((\underline{a},mp\_feedback))$
             // terminate rollout if action $\underline{a}$ is infeasible
10             if $\tau$ is $None$ then
@@ 8             $\tau,mp\_feedback\leftarrow MP(s;\underline{a})$
12
            // update state $s$ and add $\underline{a}$ to $plan$
13             $s\leftarrow\mathcal{T}(s,\underline{a}(\tau))$, $plan.append(\underline{a}(\tau))$
14
      // check whether $s$ satisfies goal
15       $success,task\_feedback\leftarrow g(s)$
       // when the plan is feasible but doesn’t reach goal
16       if not $success$ then
             // record $task\_feedback$
17             $feedback.append(task\_feedback)$
             // add aggregated feedback to $trace$
18             $trace.append(feedback)$, $trace\leftarrow trace[-k:]$
             // clear aggregated feedback
19             $feedback\leftarrow[]$
20
21
22return $success$, $plan$
@@ III-B Reasoning and Planning with pre-trained LLM
Following previous attempts in utilizing LLMs for reasoning and planning [11, 12, 13], we prompt a pre-trained LLM to generate motion failure reasoning and action sequences in text format. Since we want to limit the domain-specific prior provided to the LLM, we use zero-shot prompting without providing any planning examples. We adopt Chain-of-Thought (CoT) prompting to have the LLM generate the reasoning and action sequence in an auto-regressive manner. The prompt fed into LLM comprises the following contents: (i) a system message that provides the global context to the pre-trained LLM and activates its planning capability, (ii) a task description specifying the environment, goal, and available primitive actions, (iii) the textualized initial environment state, (iv) the trace of motion planning feedback, and (v) the output format. Note that most promp content will remain unchanged for different planning iterations of the same TAMP problem. Only the motion planning feedback trace in (iv) will be updated as new failed action sequences are added. Fig. 2(a) illustrated an example of reasoning and planning with LLM.
@@
System message: You are an AI robot that generates a plan of actions to reach the goal… You are expected to correct the plan incrementally (on top of the last plan) to avoid motion failure. This may involve sample new parameters for the failed action or reverse one or more succeeded actions for backtracking… You are expected to generate a plan from scratch.
Task description: A robot arm is tasked to pack boxes into a basket on a table. The robot sits at (0, 0), and faces the positive x-axis, while the positive z-axis points up. The robot is equipped with primitive actions, each taking a list of objects and continuous parameters as input:
$\texttt{pick}([\texttt{obj}],\{\})$: pick up obj, with no parameters. $\texttt{place}([\texttt{obj}],\{``x``:[0.0,1.0],``y``:[-1.0,1.0],\\ ``theta``:[-3.14,3.14]\})$: place obj at location $(x,y)$ with the planar rotation theta, where $x$ ranges $(0.0,1.0)$, $y$ ranges $(-1.0,1.0)$, and $theta$ ranges $(-3.14,3.14)$. Initial state: {init_state}
motion planning feedback trace: {trace}
Output format: Please generate output step-by-step, which includes your reasoning for the failure of the last plan as well as the generated plan… Answer the questions: (i) what is the cause of the failure of the last plan? (ii) can altering action parameters for the failed action solve the problem… (iii) do we need to reverse one or more succeeded actions executed before the failed action… (ii) what is your strategy to generate a new plan from scratch to accomplish the task goal? Please organize the output following the JSON format below: { “Reasoning”: “My reasoning for the failure of the last plan is …”, “Full Plan”: [“pick([’red_box’], {})”, “place([’red_box’], {’x’: 0.51, ’y’: 0.02, ’theta’: 0.00})”, …] }
@@ III-C Synthesizing Motion Planning Feedback
    The goal configuration has no feasible IK solution.
@@ IV-A Simulation Setup
The simulation has an important role in Robotics[38, 39]. We developed a PyBullet-based simulation environment for our box-packing tasks, as illustrated in Fig. 5a. In Setting 1, three different sets of objects are given with increasing total sizes, while the basket size remains unchanged. This task requires the LLM3 to oversee potential collisions among objects and the robot throughout the action sequence. The LLM must reason why collisions occur and adjust previous actions to ensure feasible task and motion plans. Setting 2 involves placing the Set 3 objects into baskets of increasing sizes. Here, the robot cannot access the entire basket region but encounters a collision likelihood similar to the most crowded condition in Setting 1. In this setting, the task becomes more complex as it challenges the LLM to reason about the robot’s operational space and adjust the plan accordingly. Throughout the simulations, we utilize GPT-4 Turbo as the LLM planner and BiRRT [37] as the motion planner, with 10 attempts for each setting.  [^1] The code is available at https://github.com/AssassinWS/LLM-TAMP.
@@ IV-B Ablation Study
    LLM3 Scratch: The proposed LLM3 framework from scratch variant. It replans the entire action sequence, incorporating feedback from the motion planner if any action fails. See Fig. 3
@@
In both settings, integrating motion planning feedback results in a decrease in #LM and #MP, along with an increase in %SR for both backtrack and scratch strategies. This indicates that the LLM can reason about failures from motion planning feedback, and importantly, propose adjusted task plans and action parameters that are more likely to produce feasible motions. Surprisingly, no clear evidence suggests that utilizing backtracking is superior to replanning from scratch. We’ve observed distinct behaviors among different strategies employed by LLM3. Specifically, when utilizing backtracking, LLM sometimes consistently adjusts the action parameter of the specific action that directly led to failure, without adjusting previous actions. In contrast, the LLM3 employing a planning from scratch strategy simply re-samples all actions and parameters, occasionally resulting in a slightly lower #LM.
@@ IV-C Action Parameter Selection
Random Samples 109.6 663.1
LLM 10.8 70.2
LLM + Feedback 7.9 53.2
@@ IV-D Experiment Setup
To validate the effectiveness of our proposed method in a real-world setting, we conducted an experiment using a Franka Research 3 manipulator. The goal was to demonstrate the robot’s ability to perceive and manipulate objects in an environment with uncertainties in both perception and execution. The robot observed a single point cloud from a third-person-view RGB-D camera, capturing the workspace containing various objects such as blocks and a plate. To identify and locate individual objects, we employed Grounded Segment Anything [40] for object segmentation, initially segmenting objects in the 2D RGB image and then projecting the results onto the corresponding 3D point cloud. This approach yielded per-object point clouds, essential for planning and executing manipulation tasks.